Find the surface area of a cube with edges 7 cm long
Answer by addingup(3677)

Area of the square: A = S^2 or you can do A = S*S which is the same thing.
A cube has 6 faces:
A(c) = 6(S*S) = 6(7*7) Use your calculator to find your answer.
`````````````````````````````````````````````````````````````
Note: Because of the Associative Property of multiplication it doesn't matter how you multiply them:
(6*7)*7 = 6(7*7)= 6*7*7